OpenStudy (anonymous):

solve the linear system using elimination x+y=2 2x +7y=9

OpenStudy (anonymous):

x + y = 2 times 2 -> 2x + 2y = 4 2x + 2y = 4 - 2x + 7y = 9 --------------------- -5y = -5 and so on

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

but its wrong

OpenStudy (anonymous):

its (1,1)